OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. x-intercept is 3, y-intercept is 2/3 The slope-intercept form for the line is?

OpenStudy (anonymous):

Can anyone help me finding the slope-intercept?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

first tell me what do you understand by x intercept?

OpenStudy (anonymous):

i mean what will be the y coordinate

OpenStudy (anonymous):

I know we have to do y=mx+b

OpenStudy (anonymous):

lets understand few things x intercept implies line crossed x axis, so y coordiante is 0 so the point has the values or coordinates as (3,0) similarly for y intercept x coordiante is 0 so the point has coordinates as (0,2/3)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now find slope m= (2/3-0)/(0-3)

OpenStudy (anonymous):

which works out to be -2/9

OpenStudy (anonymous):

now use y-y1=m(x-x1) where (x1,y1) can be either (3,0) or (0,2/3)

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

how did you get -2/9

OpenStudy (anonymous):

(2/3-0)/(0-3)

hero (hero):

So basically you have two points (3,0) and (0,2/3) Now insert each point into y = mx + b and solve for m 0 = 3m + b 2/3 = 0m + b -3m = b 2/3 = b -3m = 2/3 m = -2/9 Therefore the equation of the line is y = -2/9(m) + 2/3

OpenStudy (anonymous):

Write the function as a set of order pairs. Give the domain and range of it. f(a)=r, f(b)=p, f(c)=t, f(d)=r

OpenStudy (anonymous):

Can someone help me?